# SurgSora: Decoupled RGBD-Flow Diffusion Model for Controllable Surgical Video Generation

### v1

@@ 4.3.1 Video Generation Evaluation and Visualization
MOFA-Video [27] 95.59% 671.66 692.44 96.31 2.685
SurgSora (Ours) 98.70% 395.65 535.95 87.94 3.278
Table 1 and Table 2 quantitatively compare the performance of SurgSora with existing methods. The two unconditional video generative models, Endora [23] and StyleGAN-V [33] take random noise as the input, while the conditional video generative models, MOFA-Video [27] and our SurgSora take the first image frame as the input. Here, to facilitate the comparison with the ground-truth videos, we employ the optical flows estimated from the ground-truth video as the trajectory control.
As can be seen from Table 1, our framework exhibits the highest Frame Consistency at 98.70%, which is a modest improvement over StyleGAN-V’s 98.02% and a more pronounced leap from Endora’s 97.51%. This indicates our model’s ability to maintain uniformity and temporal stability across video frames more effectively. More importantly, in terms of video quality metrics, our method achieves the lowest FVD at 395.65 as compared to StyleGAN-V and Endora, which record FVDs of 857.16 and 1146.54, respectively. The CD-FVD, which assesses content and dynamics in video synthesis, further underscores our method’s effectiveness with a score of 535.95, markedly better than the closest method, MOFA-Video [27] at 671.66.
@@
Figure 4: Comparison of conditional generation results trained on CoPESD dataset: Generated video by ours retains most of the authenticity of the real video. The video generated by MOFA-Video [27] has a serious distortion at the end of the video.
Table 2: Quantitative Comparisons on CoPESD Dataset [38]: This table evaluates conditionally generated video and optical flow qualities.
Models Video Optical Flow
[…]
MOFA-Video [27] 19.06 48.28% 0.2620 265.54
SurgSora (Ours) 20.71 55.94% 0.1477 149.89
Furthermore, our approach achieves the best FID of 87.94, improving upon MOFA-Video’s 96.31 and significantly outperforming Endora’s 205.93. This suggests that the videos generated by SurgSora are visually closer to authentic videos, thereby capturing the nuances of video content with greater fidelity. SurgSora outperforms the IS score (3.278) among all, which indicates better-generated frames in terms of objectness and diversity.
Moreover, for conditional generation, as shown in Table 2, SurgSora outperforms MOFA-Video with scores of 20.71 dB in PSNR and 55.94% in SSIM compared to MOFA-Video (19.06 dB and 48.28%), which confirms our model’s superior capacity in preserving generation quality and structural similarity. Notably, we evaluate the model’s ability to adhere to trajectory control by comparing the given optical flow with the optical flow extracted from the generated video. Our SurgSora framework consistently demonstrates superior accuracy in following the specified generation instructions.
Visual comparisons are given in Figure 3 for unconditional generation and Figure 4 for conditional generation. Unconditionally generated videos from both Endora [23] and StyleGAN-V [33] do not conform to reality in terms of the generation of scenes and instruments, which cannot accurately reflect real medical situations. Conditionally generated video by MOFA-Video [27] involves serious frame distortion at the end of the video. Our SurgSora-generated video maintains most of the real video’s authenticity, further showing its superior performance.
Figure 5: The quantitative results of our SurgSora with different trajectory conditions.
4.3.2 Customize Trajectory Video Generation
Figure 6: The quantitative results of our SurgSora using different trajectory conditions: We calculated the difference between the first and last frames and visualized it as a heatmap, where red means high movements and blue represents low movements. Zoom in to see the details.
@@ 4.3.2 Customize Trajectory Video Generation
To address the effectiveness of our Trajectory Controller block, we generate a few demos by using the TC module. Figure 5 shows videos generated from varying surgical image trajectories, clearly depicting the dynamic movement and transformation of the objects within the images in accordance with the specified trajectories. Further demonstrating the module’s capacity for precise control, we generated distinct trajectories within the same image, as presented in Figure 6. We manipulated tissues and instruments to move in designated directions, and the visual results showed that objects along the set paths moved without noticeable distortion. The heatmap indicates that devices have undergone obvious changes according to the trajectory requirements while maintaining the background unchanged. These outcomes validate the high performance and accuracy of our module in controlling and generating detailed movement in medical imagery.
4.3.3 Ablation Study
Table 3: Ablation experiments of our SurgSora on the CoPESD Dataset [38]. To observe the performance changes, we (i) remove the Segment Feature, (ii) degenerate the depth branch, and (iii) discard the Multi-Scale Fusion Block.
@@ 4.3.3 Ablation Study
We conduct ablation studies on our SurgSora model using the CoPESD Dataset [38] to demonstrate the contribution of each component, with results summarized in Table 3. Comparing Rows 3 and 5 in the Table, disabling the segment feature while retaining the depth branch and Multi-Scale Fusion (MSF) block leads to a performance drop, with FVD/CD-FVD increasing from 395.65/535.95 to 479.13/624.63, highlighting the segment feature’s role in enhancing visual coherence and temporal consistency. Comparing Rows 2 and 3, further removing the MSF block results in a substantial decline, with Frame Consistency dropping to 96.99% and FVD/CD-FVD increasing significantly to 510.11/782.51, indicating the depth branch’s importance in integrating spatial information. Comparing Rows 4 and 5, when the MSF block alone is disabled but the segment feature and depth branch remain active, the performance degradation is less severe, with Frame Consistency at 97.53% and moderate increases in FVD/CD-FVD (422.06/603.34), showing that the MSF block enhances feature fusion but the segment feature and depth branch remain crucial. Direct use of decoupled-flow features produces the worst SSIM score (51.33%), suggesting poor structural preservation, which improves with the depth branch but still incurs computational complexity and lower performance. The addition of the MSF block resolves this issue by aligning the decoupled features with the overall framework, significantly improving metrics. In Row 5, integrating all three components—segment feature, depth branch, and MSF block—our complete SurgSora achieves the best performance, demonstrating their collective effectiveness in enhancing video quality, structural consistency, and temporal coherence.
5 Limitation
The accuracy and authenticity of videos are particularly important in the medical field. Moreover, the complexity of medical videos is much higher than that of general video content, containing complex dynamic structures and rich details. SurgSora has been evaluated on specific datasets, but its ability to generalize across surgical scenarios and different types of surgeries still needs to be questioned. Besides that, generating long clips requires the model to sustain narrative coherence and visual consistency, which remains challenging due to accumulating errors and drifts in the generated content over time. The model’s current capabilities may not fully account for the dynamic and unpredictable nature of live surgeries, where multiple instruments and varying anatomical structures interact in complex ways. This limitation affects the model’s utility in training and planning, where understanding these interactions is crucial. Addressing these limitations requires advancements in computational strategies and more robust adaptability of the model to diverse medical contexts. By overcoming these challenges, medical video generation technology can significantly improve, offering more reliable tools for medical training and procedural planning.

### v2

@@ 3.1 Experimental Results
MOFA-Video [12] 95.59% 671.66 692.44 96.31 2.685 19.06 48.28% 0.2620 265.54
SurgSora (Ours) 53.33% 98.70% 395.65 535.95 87.94 3.278 20.71 55.94% 0.1477 149.89
Figure 2: Comparison of unconditional results on CoPESD dataset.
Table 1 quantitatively assesses SurgSora alongside existing generative models, highlighting our approach’s superior performance in video generation. SurgSora achieves the highest frame consistency at 98.70%, surpassing Endora (97.51%) and StyleGAN-V (98.02%), indicating enhanced temporal stability. Besides, it also achieves the best FVD score of 395.65, outperforming StyleGAN-V and Endora, which showcase FVD scores of 857.16 and 1146.54, respectively. Additionally, SurgSora excels in content and dynamics with a CD-FVD score of 535.95, markedly better than MOFA-Video (671.66). Furthermore, SurgSora leads in visual fidelity with the best FID of 87.94, demonstrating its capacity to generate videos that closely mimic real surgical scenes. The model also tops IS at 3.278, indicating superior object and diversity representation.
@@
Figure 3: Comparison of conditional generation results trained on CoPESD dataset.
For conditional generation, SurgSora outperforms MOFA-Video with scores of 20.71 dB in PSNR and 55.94% in SSIM compared to MOFA-Video (19.06 dB and 48.28%), which confirms promising capacity in preserving generation quality and structural similarity. Visual evidence from Fig. 2 and Fig. 3 shows SurgSora’s superior performance, with videos that maintain authenticity and suffer less distortion compared to other models, showing its practical efficacy in medical applications.
@@
Customize Trajectory Video Generation. To address the effectiveness of our Trajectory Controller block, we generate a few demos by using the TC module. Fig. 4 (a)(b) shows videos generated from varying surgical image trajectories, clearly depicting the dynamic movement and transformation of the objects within the images in accordance with the specified trajectories. Further demonstrating the module’s capacity for precise control, we generated distinct trajectories within the same image, as presented in Fig. 4(c)(d)(e). We manipulated tissues and instruments to move in designated directions, and the visual results showed that objects along the set paths moved without noticeable distortion. The heatmap indicates that devices have undergone obvious changes according to the trajectory requirements while maintaining the background unchanged. These outcomes validate the high performance and accuracy of our module in controlling and generating detailed movement in medical imagery.
Human Evaluation. A blind test was conducted to assess the perceptual quality of video clips generated by SurgSora. We randomly selected seven clips from SurgSora-generated videos and eight from real surgical recordings (15 clips in total). These clips were shuffled and anonymized before being presented to six surgeons, who were individually asked to distinguish between generated and real videos. The results, shown in Fig. 1, indicate an average accuracy of 53.33%, which is close to random guessing (50%), suggesting that SurgSora produces highly realistic surgical videos that are difficult to differentiate from real footage.
Figure 4: Quantitative results of SurgSora within different trajectories and samples.
Table 2: Ablation experiments of our SurgSora on the CoPESD Dataset [21].
@@
Ablation Study. We carried out ablation studies on the SurgSora model using the CoPESD Dataset [21] to assess the impact of different components, summarized in Table 2. Results highlight that removing the segment feature increases FVD/CD-FVD from 395.65/535.95 to 479.13/624.63, underscoring its role in improving visual and temporal coherence. Eliminating the MSF block further degrades performance, with Frame Consistency dropping to 96.99% and a larger increase in FVD/CD-FVD to 510.11/782.51, emphasizing the depth branch’s role in spatial integration). Disabling only the MSF block results in a milder performance drop. The worst structural preservation indicated by the lowest SSIM occurs when decoupled-flow features are used alone, but performance improves with the addition of the depth branch. Integrating all three components yields the best results, confirming their collective utility in video generation quality.